# NPHardEval4V: A Dynamic Reasoning Benchmark of Multimodal Large Language Models

### v2

@@ Appendix B Recognition Accuracy
Figure 10: Prompt recognition across models, complexity levels, and difficulty levels
Recognition is an important preprocess that disentangles input processing’s impact on reasoning ability of MLLMs. In general, close source models outperforms open source counterparts in recognition. It is also important to notice irrelevance between complexity of tasks and recognition accuracy.
Several close source models, however, can obtain similar or even better recognition ability than the open source ones. For instance, CogVLM constantly reaches full recognition rate (1.0) on SAS and KSP problems. LLaVa, Otter, and QWen-VL also have different extents of outstanding recognition abilities on different tasks. In addition to further showing MLLMs’ variation in tasks regarding recognition ability, this result also reveals the individual strengths of different MLLMs, possibly due to different training data and techniques.

### v3

@@ Appendix C Recognition Accuracy
Figure 9. Prompt recognition across models, complexity levels, and difficulty levels
Recognition is a critical initial step, as a model’s ability to understand the prompt directly impacts its performance on downstream tasks. As shown in Figure 9, closed-source models like gpt-4o-2024-11-20, gemini-2.0-flash-exp, and claude-3-7-sonnet-20250219 generally exhibit superior and more consistent recognition accuracy across the majority of tasks. It is also notable that performance varies dramatically between different task types; for example, most models excel at TSP_Decision while struggling significantly with TSP, indicating that task formulation is a key factor in recognition success.
Despite this trend, open-source models demonstrate impressive recognition capabilities, matching or exceeding their closed-source counterparts on specific tasks. For instance, Intern achieves perfect recognition (1.0) on several difficulty levels of the BSP task. Similarly, models like ovis, nvlm, and qwen25vl72b show outstanding performance, particularly on the TSP_Decision and GCP_Decision tasks, often maintaining near-perfect accuracy. This not only highlights the significant variation in prompt recognition ability across different tasks but also underscores the unique specializations of individual LVLMs, likely stemming from differences in their training datasets and architectural designs.